OpenStudy (anonymous):

Determine between which consecutive integers the real zeros of each function are located. f(x)= 2x^2 - 5x + 1

OpenStudy (tkhunny):

f(0) = 1 f(1) = 2 - 5 + 1 = -2 [There is a Real Zero between 0 and 1] f(2) = 8 - 10 + 1 = -1 f(3) = 18 - 15 + 1 = 4 [There is a Real Zero between 2 and 3] Really, that's all it takes.

OpenStudy (anonymous):

Okay thanks! I was totally making that harder than it is.